### v1

# On The Topology of Polygonal Meshes

Abstract
This paper is an introductory and informal exposition on the topology of polygonal meshes. We begin with a broad overview of topological notions and discuss how homeomorphisms, homotopy, and homology can be used to characterize topology. We move on to define polygonal meshes and make a distinction between intrinsic topology and extrinsic topology which depends on the embedding space. A distinction is also made between quantitative topological properties and qualitative properties. We outline a proof of the Euler and the Euler-Poincaré formulas. The Betti numbers are defined in terms of the Euler-Poincaré formula and other mesh statistics rather than as cardinalities of the homology groups which allows us to avoid abstract algebra. Finally, we discuss how it is possible to cut a polygonal mesh such that it becomes a topological disc.
1 Introduction
Figure 1: The Fertility model is shown on the left. It may not be immediately obvious how many holes (or handles) it possesses, but on the right we see a sphere with four handles which is topologically equivalent, i.e. there is a homeomorphism between these two shapes, and on the right it is obvious that the model has four handles.
Topology is the study of invariants of shapes under deformation. These invariants are quantities that do not change no matter how much you deform the shape – as long as it is neither torn nor glued. A specific invariant that we often think about is the number of holes, and a classical example is that a cup with a single handle is topologically equivalent to a donut.
From a practical point of view, topological properties can be of great interest. For instance, we may want to count the number of objects in a 2D or 3D image; this is such a common task that it is barely recognized as being related to topology. In other cases we may want to determine whether objects contain cycles or have a purely branching structure, or we may be interested in knowing whether an object has internal cavities.
In some situations we may be interested in topological properties because they have some biological significance [Arnavaz2022]. In other situations we may need to compute topological properties in order to define loss functions for machine learning [clough2020topological]. Finally, there are situations where we need to restrict the topology and want to detect potential topological changes in order to avoid them [Cui2024].
While there are many applications of topology, it can be a challenging topic unless one is in possession of a good mathematical background; in particular, a firm grasp of abstract algebra is needed for algebraic topology. In this document, we target applications in visual computing and sidestep some of the mathematical requirements by considering only the topology of 3D surfaces provided as polygonal meshes. Moreover, we derive all quantitative topological properties from the Euler-Poincaré formula. By limiting the scope in this way, we are able to arrive at formulas for key topological properties such as genus and Betti numbers without group theory [alexandroff2012introduction]. Since Betti numbers are almost always defined in algebraic terms this document takes an unusual route which I hope is illuminating for readers who wish to get a grasp on topology without necessarily understanding homology.
For readers who do want to understand homology, the book by Edelsbrunner and Harer [edelsbrunner2010computational] is a good introduction that focuses on the computational aspects, and Giblin’s book [giblin2013graphs] is recommended as a more mathematical text. Readers interested in a deeper understanding of the Euler-Poincaré formula should consider Imre Lakatos’ extraordinary book “Proofs and Refutations” [lakatos1963proofs] which uses the formula merely as an example, but since it is the only example it is covered very deeply.
1.1 What is a Hole?
Topological properties often feel familiar and appeal to our intuition, but subtleties arise when we try to make the notions precise. Perhaps the most pernicious source of confusion is the fact that solid objects and surfaces are very different from a topological point of view. This is true even in the cases where the surface is exactly identical to the boundary of some solid. For instance, consider an egg-like object which is hollow inside. We might consider this object to consist of two surfaces: the outer surface and the surface that encloses the interior void. However, if we restrict our topological analysis to surfaces only, there is no way to distinguish an object with an interior cavity from two objects where one just happens to be inside the other. Since we will only be dealing with polygonal meshes in the following, it is worthwhile to be quite clear that polygonal meshes are surfaces, and there is no way of telling (from the representation itself) whether two disjoint polygonal meshes are part of the boundary of the same solid.
Another sticking point is related to the very notion of a hole. Consider a piece of paper. It is obvious that we can cut a hole in the paper with a pair of scissors. Now, consider a donut. Clearly, the donut also has a hole, but it is of a different type. Holes of the donut variety are shown in Figure 1.
1.2 Background
Figure 2: Top left: A homeomorphism, $f:A\rightarrow B$, is a continuous, bijective map which has a continuous inverse. Middle bottom: A homotopy is a continuous deformation. In this example an annulus is homotopic to a ring since the former can be continuously shrunk into the latter. Top right: The red curves are homologous because they differ only by the boundary of a region. However, the red curves are not homologous to the green or blue curves. The blue curve, on the other hand, is homologous to zero because it is the boundary of a region.
[…]
Homeomorphisms
Two topological spaces are said to be homeomorphic if there is a continuous bijective map from one to the other and this map has a continuous inverse (cf. Figure 2 top left). This is the strictest test of topological equivalence, but given two objects of identical, topology there is no general method for computing a homeomorphism between them. Moreover, homeomorphisms only allow us to make statements about topological equivalence, they do not allow for quantification of topological properties. Nevertheless, we will often need to make statements about topological equivalence in terms of homeomorphisms in what follows. For instance, if we say that something is a topological disc, it means that there is a homeomorphism between the object and a disc.
Manifolds
margin: [Uncaptioned image]
A topological space, $X$, is a 2-manifold if any sufficiently small neighborhood of a point, $\mathbf{p}\in X$, is homeomorphic to an open disc (see the margin figure on the right). Intuitively, this means that we can cut out a (possibly tiny) patch which contains $\mathbf{p}$ and which can be flattened and stretched to a perfect disc. Thus, a 2-manifold models $\mathbb{R}^{2}$ locally at every point and can be seen as a mathematical definition of the notion of a closed surface. Of course, some surfaces, e.g. a sheet of paper, are open, and an open surface is a 2-manifold with boundary. Points on the boundary are homeomorphic to half-open disc as shown on the right.
The notion of a 2-manifold generalizes beyond surfaces in 3D. An $n$-manifold in $\mathbb{R}^{d}$ where $n\leq d$ is a set of points that is locally homeomorphic to either an open $n$-ball or a half-open $n$-ball in a sufficiently small region around every point belonging to the $n$-manifold. Because we will only consider surfaces, manifolds will always refer to 2-manifolds in the following.
Homotopy
Another way to characterize topology is through homotopy. Two topological spaces are considered homotopic if one can be continuously deformed into the other (cf. Figure 2 middle bottom). At first glance, it may appear to be the same thing for two objects to be homeomorphic and homotopic, but that is actually false. A solid sphere is homotopic to a single point since we can smoothly contract the sphere to a point, but there is clearly no way to establish a bijective continuous mapping with a continuous inverse from a solid sphere to a single point. Conversely, we can cut a torus, tie it into a knot, and then stitch the ends back up. The knotted torus will still be homeomorphic to the original, but it is not homotopic since the continuous deformation would get stuck on the knot.
Simplicial Homology
The notion of homology from the field of algebraic topology provides us with a tool for quantifying topological properties. We will briefly outline the ideas of simplicial homology which assumes that the object is a discrete structure which we can think of as a triangle mesh. Homology allows us to compute the so-called Betti numbers, which are a mathematically precise way to count cycles (c.f. Section 5.1) according to the dimension of the cycle, are defined as the cardinalities of the homology groups, $H_{k}$, and computed according to a deceptively simple formula,
$\beta_{k}=|H_{k}|=\left|\frac{Z_{k}}{B_{k}}\right|,$ (1)
where $Z_{k}$ is the group of all k-dimensional cycles, and $B_{k}$ are the k-dimensional boundaries. For instance, $Z_{1}$ consists of all closed curves (i.e. 1D cycles) on the shape under consideration. Likewise, $B_{1}$ is the group of all boundaries of 2-dimensional regions of the shape.
The involved entities are groups [alexandroff2012introduction]. A group is defined as a set of mathematical objects endowed with an operator that combines two objects, producing a new object which must again belong to the set, i.e. fulfilling the closure property. Perhaps the most familiar example of a group is the set of integers with addition as the operation, but collections of mesh entities called chains also constitute groups. A chain can be understood as a vector of mesh entities that can be either vertices, edges, or faces. The operator that combines two chains simply adds the entities modulo two. Hence, an entity is in the resulting chain if it belongs to either but not both of the chains combined by the operator.
Given an appropriate equivalence relation, groups can be partitioned into equivalence classes called quotient groups. If we consider, $H_{1}$, the quotient group (formed by the division-like notation in Eq. (1)) should be understood as a partitioning of all closed chains of edges ($Z_{1}$) by the equivalence relation that they differ only by the boundary of a region ($B_{1}$). In the case of an object that has sphere topology, all closed chains of edges are boundaries of some region. This means that they all differ from the empty cycle by the boundary of a region and we say that such chains are homologous to zero. On a torus there are two distinct families of chains of edges that do not bound any regions. Within these families, each chain differs from every other chain belonging to the same family by the boundary of a region, but unlike on the sphere we cannot make them go away by subtracting the boundary of a region. Summing up, on topological spheres all closed chains of edges are homologous to zero, but on topological tori there are two families of chains of edges which are not homologous to zero, and this is why $|H_{1}|=2$ for the torus. See Figure 2 (top right) where the red and green curves are representatives from each of these two families.
This brief discussion was meant to provide only the gist of homology groups. A much more thorough discussion is provided by Edelsbrunner and Harer [edelsbrunner2010computational] (and many others) but is outside the scope of this text which is rather to demonstrate how the Betti numbers can be computed without making any reference to homology. Instead, we will demonstrate that an equivalent formulation of the Betti numbers can be derived from the Euler-Poincaré formula.
2 Polygonal Meshes
A polygonal mesh is a tuple, $\mathcal{M}$ = $\langle\mathcal{V},\mathcal{E},\mathcal{F}\rangle$, where $\mathcal{V}$ is a set of vertices, $\mathcal{E}$ is a set of edges, and $\mathcal{F}$ is a set of faces. We will often need to refer to the number of vertices, $V=|\mathcal{V}|$, the number of edges, $E=|\mathcal{E}]$, and the number of faces, $F=|\mathcal{F}|$. Formally, we will define an edge, $e=\{u,v\}$ where $u,v\in\mathcal{V}$ and $e\in\mathcal{E}$, as a set of two vertices, and a face, $f=\{e_{0},...,e_{n}\}$ where $e_{i}\in\mathcal{E}$ and $f\in\mathcal{F}$, as a set of edges. By this definition, vertices that belong to an edge must also be in the set of vertices and likewise edges in a face must belong to the set of edges. On the other hand, we may allow “dangling” edges not associated with any face or loose vertices that do not belong to an edge, but in practice such unused entities are often considered defects.
For the mesh to be geometrically meaningful, there must be a point, $\mathbf{p}_{v}\in\mathbb{R}^{3}$, associated with each vertex, $v\in\mathcal{V}$. If this association has been made, we can think of vertices as points in 3D, of edges as line segments, and of faces as polygons.
Validity of Faces
It must be true that if a vertex, $v$, belongs to any edge, $e$, of a face, $f$, then it also belongs to precisely one other edge since that ensures that the edges form a simple, closed cycle. In other words, the set of edges that contain a given vertex while belonging to a given face is either empty or of cardinality two:
$\forall v\in\mathcal{V},\forall f\in\mathcal{F}:\left|\left\{e|v\in e\wedge e\in f\right\}\right|\in\{0,2\}\,.$ (2)
We must avoid faces whose boundary consists of more than a single cycle of edges. This could be a practical issue if we, say, want to make a tunnel through a cube. The issue is that topologically we cannot always distinguish this configuration from two separate meshes (e.g. cube and tunnel). Thus, it is nearly the same issue as when we discussed interior voids, but in this case it is fixable since we can add edges to the mesh connecting the tunnel to the cube.
Validity of Edges
We require that edges belong to either one or two faces, i.e.
$\forall e\in\mathcal{E}:\left|\{f|e\in f\wedge f\in\mathcal{F}\}\right|\in\{1,2\}\,.$ (3)
This requirement is part of ensuring that the mesh is manifold. The edges that belong to only one face are called boundary edges. Just like the edges that bound a face, boundary edges also form cycles which bound holes in the polygonal mesh. The fact that meshes may have boundary cycles means that polygonal meshes do not always represent solids. However, even polygonal meshes which have no boundary cycles (i.e. all edges are adjacent to exactly two faces) are usually considered to be empty shells rather than solids for topological purposes.
If manifoldness (cf. Section 2.3.1) is not a concern, this requirement can be omitted.
Validity of Vertices
Another manifoldness condition (cf. Section 2.3.1) pertains to vertices. The faces incident on a vertex must be connected via shared edges. Again, this can be omitted if manifoldness is not a concern.
Multiple Incidence
[…]
A special case that bears considering is that of multiple incidence. An edge might be incident on the same face twice as shown in the figure on the right. Likewise, a vertex can be incident on the same face multiple times. Although multiple incidence is not problematic in itself, it can lead to problems if we manipulate a mesh in a computer. For instance, we might merge the two faces that are incident on a given edge. If the two faces are really the same, we create two disconnected cycles that bound the same face. As discussed above, this should be avoided.
2.1 Computer Representation
From the perspective of computer representation, meshes can be stored in a number of ways that differ mostly in which of the three primitive types that is considered to be the primary one.
Meshes can be stored as lists of faces where each face is represented as a cyclic list of vertices; two consecutive vertices then define an edge. This representation is preferred in computer graphics for rendering, and its simplicity makes it appealing also for disc storage of meshes. However, for manipulation of meshes, an edge-based data structure is often convenient since this type of representation makes certain editing operations efficient. For most edge-based data structures the requirement that an edge can be incident on either one or two faces is a built-in constraint.
In the present context, the representation generally does not matter, but there is one further concern that should be addressed. While at least one popular data format stores the vertices directly as 3D points in space this is not a good idea if we are interested in topology. Vertices should be identified by indices, allowing us to easily find out if the corners of two polygons correspond to the same vertex (simply by checking whether the vertex indices are the same). Of course, if we are interested in the geometry of the mesh, we would then need a separate table that maps vertex indices to points in space.
One type of non-manifoldness is generally accepted in practice: sometimes, a vertex is incident on more than two boundary edges. An example is shown in the figure on the right where two triangles share a vertex. margin: [Uncaptioned image] The result is that the vertex is incident on four boundary edges. As long as the mesh representation keeps track of which boundary edges that are adjacent in the incident boundary cycles, it is not a problem to have any even number of boundary edges incident on a vertex, but it is still required that the faces incident on the vertex all come in connected components that contribute two boundary edges. Put differently, the configuration is considered valid if it can be made manifold by closing the holes.
Triangle meshes are, of course, a special type of polygonal meshes where all faces have three edges. Numerous algorithms exist that operate only on triangle meshes and not general polygonal meshes. This is not a big limitation in practice since it is always possible to triangulate the faces of a polygonal mesh, thereby obtaining a triangle mesh. Moreover, data structures for triangle meshes are simpler than those for generic polygonal meshes.
2.2 Intrinsic and Extrinsic Topology
The polygonal mesh representation encodes both the geometry and topology of shapes. While that is powerful, it can also be slightly confusing because we can consider the intrinsic topology of a polygonal mesh or its extrinsic topology of the mesh as a set of points in an embedding space. This leads to two different points of view, and it is necessary to choose a point of view before analysing the topology of a mesh.
The Extrinsic View
Extrinsically, a polygon is a set of points in space, $\bigcup_{f}P_{f}$, where $P_{f}$ is the set of points that belong to the polygonal face $f\in\mathcal{F}$. Thus, $P_{f}$ is only defined if we know the positions of the corners of $f$ and if we know how the geometry of the polygon is defined from those corners. In practice, polygons are often assumed to be almost flat and the question of how precisely the geometry is defined is ignored. As long as we stick to triangle meshes (which is often the case) this is unproblematic since a triangle is always flat. According to the extrinsic view, the polygonal mesh representation can be understood as a generalization of the (ancient) notion of a polyhedron. For instance, a cube is a polygonal mesh with eight vertices, twelve edges, and six faces. However, polyhedra usually have flat faces which is not required for polygonal meshes. We could twist the top face of the cube (relative to the bottom) and that would make the side faces non-planar. Most geometers would no longer consider it a polyhedron, but it is still a perfectly valid polygonal mesh. We conclude that a polyhedron is a polygonal mesh, but the converse is not necessarily true.
The Intrinsic View
From the intrinsic point of view, the positions of the vertices do not matter since the geometry is not important; a mesh is simply a schematic diagram of the structure of a surface. We can use this schematic to realise the mesh in 3D by assigning positions to the vertices. Edges are then typically realised as line segments connecting the vertices, and the faces become polygonal patches. However, the intrinsic topology is independent of the realisation. Expressed differently, we can see a polygonal mesh as a graph embedded in an abstract surface. Construed as a graph, some cycles of edges bound a patch on the manifold, and the faces of the polygonal mesh identify the cycles that bound patches.
In this document, we are most concerned with the intrinsic topology. Nevertheless, it is important to be aware of this distinction since the extrinsic and the intrinsic topology do not have to agree. Specifically, $\mathcal{M}$ may intersect itself in the embedding space which would often cause its extrinsic topology to differ from the intrinsic.
Some mesh configurations are valid if we are only interested in the intrinsic topology of abstract meshes, but make no sense when we consider a polygonal mesh with vertex positions in 3D and straight edges. For instance, from an intrinsic point of view, it makes sense that a vertex is connected to itself, but if we are interested in the extrinsic topology it is problematic since such an edge has no geometric meaning. Two vertices can also be connected by more than one edge. Extrinsically, this often makes the mesh non-manifold, but intrinsically it does not.
2.3 Qualitative Topology
The three topological invariants
  • manifoldness,
  • watertightness, and
  • orientability
are often important in the context of polygonal meshes. We can tell that they are topological properties since they are unaffected by a deformation of the mesh, but unlike the number of holes they are either true of false for a given mesh and thus qualitative rather than quantitative.
2.3.1 Manifoldness
Any point on a manifold has a small neighbourhood that is homeomorphic to a disc. This is trivially true for points inside faces since faces have disc topology by construction. To test a mesh, $\mathcal{M}$, for intrinsic manifoldness, we need to check that:
  • Every edge of $\mathcal{M}$ is adjacent to either one or two faces, i.e. Eq. (3) is fulfilled.
  • The faces incident on a vertex form only a single connected component. We consider faces connected if they share an edge.
Another way of stating the vertex condition is that the union of the faces incident on the vertex should remain a single connected set if the vertex were removed. If $\mathcal{M}$ passes both of these tests, we know that it can everywhere be flattened to a disc and is thus intrinsically manifold.
Provided a mesh is manifold we can also prove that boundary edges always form cycles. As stated above, the faces incident on a vertex must form a single component. This means that a vertex is always incident on either two or zero boundary edges since the incident faces must form either a fan or a disc around the vertex. margin: [Uncaptioned image] All vertices which belong to boundaries are obviously incident on two boundary edges (see the figure on the right). This means that we can alway jump from a boundary edge to the next edge across one of its vertices, but, assuming the mesh is finite, we must eventually revisit the edge we started from.
A mesh which is produced by joining two pyramids (one upside down) such that the tips of the pyramids are a shared vertex is not manifold. We could try to flatten a region around the shared vertex, but it would always end up as two discs connected at a single point. Likewise, if two cubes are placed side by side such that two edges touch (forming a shared edge) that is also a non-manifold configuration since four faces then meet at the edge where the cubes are joined. Of course, we are here assuming that the shared vertex or edge is actually shared in the representation of the polygonal mesh. Otherwise, $\mathcal{M}$ remains manifold in the intrinsic sense though not in the embedding. In practice, manifoldness is a property which can easily be true intrinsically but not extrinsically due to self-intersections. The two non-manifold configurations are shown in Figure 3.
Figure 3: Two non-manifold configurations: on the left the two tetrahedra share a vertex, and on the right, two boxes share an edge. From the images we cannot tell whether the meshes are intrinsically or extrinsically non-manifold.
2.3.2 Watertightness
margin: [Uncaptioned image]
$\mathcal{M}$ is watertight if it divides space into an exterior region, an interior region, and $\mathcal{M}$ itself. If this is true, any path from a point in the interior to a point in the exterior contains at least one point that belongs to $\mathcal{M}$. A cube (and polyhedra in general) is watertight, but remove one of its faces (as shown in the figure on the right), and the property is lost. Again, this is the extrinsic definition of watertightness. Intrinsically, $\mathcal{M}$ is watertight if it is closed, i.e. it has no boundary. A mesh which is not watertight has at least one cycle of boundary edges.
2.3.3 Orientability
$\mathcal{M}$ is orientable if we can assign consistent orientations to all its faces. Even though the faces are not necessarily flat, a face does have a top side and a bottom side, but which is which is completely arbitrary. $\mathcal{M}$ is orientable if we can assign top sides consistently.
[…]
Our procedure for setting the orientation of a face is extrinsic because it depends on choosing a direction in the embedding space. However, the cyclic ordering of the edges of a face is an intrinsic property, and whether a consistent ordering of all faces can be achieved is likewise intrinsic.
There are several famous examples of non-orientable surfaces such as the Möbius band and the Klein bottle, but in the following we will generally assume that the polygonal meshes we consider are orientable. Non-orientable surfaces are confusing in the context of the topology of polygonal meshes. It is relatively clear that a Möbius band is not homeomorphic to an annulus, so our strongest tool can distinguish this non-orientable surface from its closest orientable relative, but they have the same homotopy type, and the quantitative methods that we will discuss later cannot make this distinction: we get the same numbers for an annulus and a Möbius band. On the other hand, it is easy to detect that a surface is non-orientable. We simply have to assign consistent orientations to all its faces and realize that this is impossible. In fact, some polygonal mesh representations (in the computer science sense) cannot represent non-orientable meshes, and a seam will be introduced somewhere.
3 Euler’s Formula
Our investigation into how we can quantify topological properties of polygonal meshes begins with Euler’s formula,
VE+F=2,V-E+F=2\,, (4)
which holds for any polygonal mesh that is a topological sphere. A polygonal mesh is considered to be a topological sphere if it is homeomorphic to a sphere. Some of the most famous examples of polygonal meshes which are also topological spheres are the Platonic solids which are shown in Figure 4. By consulting Table 1 we can ascertain that they all fulfil Eq. 4.
Figure 4: The Platonic solids are shown above. From left to right: the tetrahedron, cube, octahedron, dodecahedron, and icosahedron.
margin: [Uncaptioned image] This mesh is a topological sphere.
Of course, the Platonic solids are special because they have perfect symmetry in the sense that all faces and all vertices look exactly the same. This is not true of the mesh shown on the right. Yet the version of the Stanford Bunny that is shown in the margin is also a topological sphere, and if we plug the numbers from Table 1 into Eq. (4) it holds.
Table 1: Vertex ($V$), edge ($E$) and face ($F$) counts for the Platonic solids and the Stanford bunny. All examples have Euler characteristic VE+F=2V-E+F=2.
Name $V$ $E$ $F$
Tetrahedron 4 6 4
Cube 8 12 6
Octahedron 6 12 8
Icosahedron 20 30 12
Dodecahedron 12 30 20
Stanford Bunny 3485 10449 6966
[…]
Figure 5: A simplified version of the Stanford Armadillo is shown left. On the right is a mesh consisting of the exact same vertices and faces but transformed such that all vertices lie on the sphere. The edges remain straight lines, but it is east to imagine that they could be replaced with great arcs: if we imagine a light placed in the centre of the sphere, the shadows that the edges would cast on the (inside of) the sphere would be precisely the needed circular arcs as explained by Eppstein in his Geometry Junkyard.
We begin by observing that the area of the (unit) sphere is $4\pi$. Thus, since the spherical polygons obtained from mapping onto the sphere must cover the sphere precisely, it is true that
$\sum_{f=1}^{\mathcal{F}}A_{f}=4\pi\,$, (5)
where $A_{f}$ is the spherical area of face $f$.
We know that the area of a spherical polygon is given by Harriot’s formula,
$A_{f}=\sum_{i=1}^{k_{f}}\alpha_{i}-(k_{f}-2)\pi\,$, (6)
where $k_{f}$ is the number of sides of $f$ and $\alpha_{i}$ is the angle of corner $i$ of the polygon. See Tristan Needham [needham2021visual] for a proof of this formula for spherical triangles. The generalization to spherical polygons is straight forward since polygons can always be triangulated. We can plug (6) into (5), obtaining
$\sum_{f=1}^{\mathcal{F}}\left(\sum_{i=1}^{k_{f}}\alpha_{i}-(k_{f}-2)\pi\right)=4\pi$
To proceed, we use the fact that the angles around each vertex sum to $2\pi$, though it may be worth pointing out that this is only true because the polygons have been transformed into spherical polygons. If the edges had been straight then the angle sum would be less than $2\pi$ since the corners would not have been flat in the limit. Now, we have
$2\pi V-\sum_{f=1}^{\mathcal{F}}(k_{f}-2)\pi=4\pi$
The final observation that we require is that the sum of the sides of the faces, $k_{f}$, is twice the number of edges, $E$, since every edge is shared by two faces. Reducing the formula, we get
$2\pi V-2\pi E+2\pi F=4\pi$
from which we can obtain Euler’s formula just by dividing by $2\pi$.
4 Quantitative Topology
Some topological properties are quantitative. In the context of polygonal meshes or surface more generally, we identify the numbers of
  • shells (or connected components),
  • holes (referred to as genus), and
  • boundary curves (which, confusingly, are also thought of as holes).
The number of shells is simply the number of separate pieces of the mesh, and the number of boundary curves is the number of distinct boundary cycles. Finally, the genus is the number of holes (in the sense of a hole in a donut) also sometimes referred to as the number of handles of the shape since any hole can be transformed into a handle without tearing or stitching the shape which means that holes and handles are the same from a topological point of view. Figure 1 shows two objects both of genus four, but on the right it is much more obvious that the genus corresponds to the number of handles than on the left.
None of these three properties can be obtained directly from Euler’s formula. Fortunately, there is a generalization known as the Euler-Poincaré formula:
$V-E+F=2(s-g)-b=\chi\,$, (7)
where $s$ is the number of connected components (or shells), $g$ is the genus, $b$ is the number of boundary curves, and $\chi$ is known as the Euler characteristic.
Shells
$s$ is easy to interpret. It is simply the number of connected components of the mesh. As, we recall there is no notion of solids when we deal with polygonal meshes. Hence, we do not distinguish between mesh components that represent interior voids and separate shapes even if it may be easy to tell these cases apart based on the geometry. If the mesh does have sphere topology, $s=1$, $g=0$, and $b=0$, and we immediately see that the formula reduces to the familiar one in this case. The generalization to $s>1$ is trivial. If we have two objects of sphere topology and add up their vertices, edges, and faces, we get
VE+F=2s.V-E+F=2s\,.
Proceeding to polygonal meshes with boundaries, we can assume that there is only a single component, $s=1$, without loss of generality.
Boundary Curves
$b$ is the number of boundary curves. Let us first consider the simple case where a boundary is introduced by removing just a single face. It is clear that
$V-E+F^{\prime}=2-1$
where $F^{\prime}=F-1$ since we removed one face. In other words, removing a single face just amounts to subtracting one from the right-hand side. What if the hole is bigger than a single face? We soon realise that removing more than one face does not change the left-hand side: when one more faces is removed to enlarge a hole, we also remove the edge that separates this faces from the hole. In some cases a face shares more than one edge with a hole, but then a vertex is also removed, and the end result would always be that the left-hand side is not changed further. We arrive at
VE+F=2sbV-E+F=2s-b
Genus
Arguably, $g$, the genus of the shape, is the “poster entity” of topological invariants, and it is equal to the number of handles of the shape. A handle is the same as a hole, since we can always deform (and homeomorphically map) a shape with $N$ holes into a sphere with $N$ handles. Thus, $g=1$ for a torus which has a single hole/handle and $g=0$ for a sphere that has none.
To understand how $g$ affects the Euler formula, we only have to consider what happens when we cut a handle [courant1996mathematics]. The handle is part of the mesh, and cutting it we would cut $N$ edges and slice through their incident $N$ faces. Where we cut each edge we need to insert two new vertices – one on either side of the cut – and the new vertices need to be connected by edges – again an edge on either side. This means that we introduce $2N$ new vertices and $3N$ new edges since we split $N$ edges and need $2N$ edges to connect the new vertices. We also split $N$ faces introducing $N$ new faces. Putting this together we see that
$V^{\prime}-E^{\prime}+F^{\prime}=(V+2N)-(E+3N)+(F+N)=V-E+F\,.$
Yet something changed since we never closed the two holes where the handle was cut as shown in Figure 6. Thus, in terms of the Euler characteristic, a handle is equivalent to two holes. It must haven been true that VE+F=22=0V-E+F=2-2=0 before we made the cut, if $\mathcal{M}$ was a topological torus with a single handle. It follows that we have to reduce the right-hand side by two for every handle. This leads to
VE+F=2s2gb.V-E+F=2s-2g-b\,.
Thus, the general Euler-Poincaré formula has been proven. To be fair, our proof does not meet the standards of rigor you would expect from a text aimed at mathematicians, but leaving that aside we can be certain that Eq. (7) holds.
Figure 6: Left: A polygonal mesh with handle. Right: The same mesh after the handle was cut. It is easy to see how the geometry has proliferated due to the cut. There is an additional eight vertices, twelve edges and four faces. Moreover, the cut has introduced two holes in the mesh.
4.0.1 Computing $s$, $g$, and $b$
Fortunately, it is easy to find the values of $s$, $g$, and $b$ if we are given a polygonal mesh, $\mathcal{M}$. Finding $V$, $E$, and $F$ on the left hand side of (7) is just a matter of counting. However, we must remember to only count edges and vertices once when they are shared by several faces of $\mathcal{M}$. If the faces of $\mathcal{M}$ are stored in such a way that we cannot tell whether edges or vertices are shared, it is necessary to first match up corners and polygon edges that correspond to identical vertices and edges of $\mathcal{M}$.
$s$ can be found using “flood filling”. Start from any given vertex and visit adjacent, unvisited vertices recursively. When this procedure terminates, pick an unvisited vertex and repeat. $s$ is the number of times this process is started.
$b$ can be found in a very similar fashion. We can identify boundary edges because they only have a single incident face. Starting from any boundary edge, we visit all the connected boundary edges, tracing out its cycle (c.f. Section 5.1). Once a boundary cycle has been found, we mark all its edges as visited and pick the next unvisited edge. $b$ is the number of times this process was started.
$g$ is now the only term that is missing and we can easily solve for $g$ using (7).
5 The Betti Numbers
Having proven that the Euler-Poincaré formula is true does not seem to have been very helpful for understanding it. However, there is a clue in its structure: it looks like the beginning of an alternating sum since the first term, $V$, is added, the next term, $E$, is subtracted, and the final term, $F$, is again added. We will now show that VE+FV-E+F is equal to another alternating sum which has a much more clear interpretation. To do so, we need to make the notion of cycles precise.
5.1 Cycles
Cycles are sets of mesh entities which have no boundary. While the name evokes “cycles of edges”, we will also define vertex cycles and face cycles.
Vertex Cycles
Any single vertex is a cycle. This is simply a matter of definition: we define a vertex to have no boundary. Consequently, a vertex is a cycle. Any collection of vertices is also a cycle.
Edge Cycles
Loosely speaking, an edge cycle is a set of edges that “bites its own tail”. Eq. (2) lays down the more precise condition that the edges of a face form a cycle if any vertex of the mesh belongs to either zero or two edges of the face. Note that we can make more complex cycles by combining such simple cycles and in this case a vertex can be included more than two times, but always an even number of times.
Face Cycles
A set of faces is a cycle if every edge of every face is adjacent to precisely two faces. In other words, a set of faces is a cycle if it is closed (or watertight). Again, a collection of face cycles is still a cycle.
Please refer to Figure 7 for an illustration of edge cycles and face cycles.
Figure 7: Left: two sets of edges where the left one is not a cycle, but the right one is a cycle. The cycle instigator is shown as a red line segment. Right: a cube which is not a cycle since it is missing a face is shown on the left. On the right a closed cube which constitutes a cycle.
5.2 Cycle Instigators
The next task is to count how many vertices, edges, and faces, we can choose without creating any cycles. To this end, we will introduce the notion of a cycle instigator. Given a set of mesh edges, a cycle instigator is an edge, $e^{i}$, that becomes part of a cycle when added to the set. To be precise, $e^{i}$ is a cycle instigator with respect to a set of edges $\mathcal{T}$ iff
$\exists\mathcal{C}:\mathcal{C}\subseteq\mathcal{T}\cup e^{i}\wedge e^{i}\in\mathcal{C}$
where $\mathcal{C}$ is a cycle. Now, the question we ask is how many edges from the total set of edges, $\mathcal{E}$, we can add to $\mathcal{T}$ without breaking the property that $\mathcal{T}$ contains no cycles? Put differently, what is the maximal number of edges we can add without including an instigator?
Figure 8: A tree is shown on the left with a black dot that indicates the root. In the middle image, a cycle-instigator has been added (red), and on the right the orange and green curves connect the end points of the instigator edge with the root. Adjoining the instigator edge with the orange and green edges, we clearly form a cycle (that bounds the pink region).
In fact, this question is easy to answer. Assuming $\mathcal{M}$ is a single component, a spanning tree of the vertices of $\mathcal{M}$ is a collection of edges which is incident on every vertex of $\mathcal{V}$. Now, a tree contains no cycles by definition, and we can create a tree by first adding any edge of the mesh to $\mathcal{T}$ and subsequently adding only edges that connect a vertex already incident on $\mathcal{T}$ to a vertex on which $\mathcal{T}$ is not incident. If we count the edges while adding them, it is clear that our count reaches V1V-1 since we add one edge for every vertex except the first. To answer the question, we now note that any edge that does not belong to the tree defines a unique cycle and is thus an instigator. We can obtain this cycle by adjoining the instigating edge to the path we obtain by following the tree edges from its end points back towards the root until they meet at a vertex (which may be the root vertex). This procedure is shown in Figure 8. We conclude that the maximal number of non-instigating edges is the number of edges in a spanning tree, i.e. V1V-1, and we can now define the number of non-instigating edges, $E_{N}$, and the number of cycle-instigating edges, $E_{C}$, as follows:
$E_{N}$ $=$ $V-1\,$, (8)
$E_{C}$ $=$ $E-E_{N}=E-V+1\,.$ (9)
Note that in the case of multiple connected components
$E_{N}$ $=$ $V-s\,$, (10)
$E_{C}$ $=$ $E-E_{N}=E-V+s\,.$ (11)
since we have to construct a tree for each component, $i$, and the tree for each component consists of $V_{i}-1$ edges where $\sum_{i\in[1..s]}V_{i}=V$.
In the case of vertices, the task of dividing up the vertices is trivial. All vertices are cycles, hence all vertices instigate cycles, and
$V_{N}$ $=$ $0\,$, (12)
$V_{C}$ $=$ $V\,.$ (13)
For faces, the situation is also straightforward. A mesh is a cycle only if it is closed. This means that we can pick all but one of the faces without creating a cycle. In other words, if the mesh is watertight, the maximum (and minimum) number of non-instigating faces is the number of faces minus 1:
$F_{N}$ $=$ $F-1\,$, (14)
$F_{C}$ $=$ $1$ (15)
If the mesh is not watertight, $F_{N}=F$ and $F_{C}=0$. When we are dealing with meshes that consist of multiple components which may or may not be watertight, then
$F_{N}$ $=$ $F-s_{w}\,$, (16)
$F_{C}$ $=$ $s_{w}\,$, (17)
where $s_{w}$ is the number of watertight components.
In all cases, the number of instigators and non-instigators sum to the original number of entities: $V=V_{N}+V_{C}$, $E=E_{N}+E_{C}$, and $F=F_{N}+F_{C}$.
5.3 Reformulating Euler’s Formula
With the division of mesh entities into instigators and non-instigators of cycles, we can express the right-hand side of the Euler-Poincaré formula slightly differently,
$V-E+F=(V_{N}+V_{C})-(E_{N}+E_{C})+(F_{N}+F_{C})\,$,
and rearrange the terms by shifting the non-instigating terms to the left,
$V-E+F=(V_{C}-E_{N})-(E_{C}-F_{N})+(F_{C}-0)\,$,
since $V_{N}=0$ and there is no non-instigating term of higher dimension than that for faces. Let us rename the three terms on the right-hand side:
$\beta_{0}$ $=$ $V_{C}-E_{N}\,$, (18)
$\beta_{1}$ $=$ $E_{C}-F_{N}\,$, (19)
$\beta_{2}$ $=$ $F_{C}-0\,.$ (20)
We now obtain
$V-E+F=\beta_{0}-\beta_{1}+\beta_{2}\,.$ (21)
The left-hand side of (21) is an alternating sum of mesh elements that increase in dimension. The right-hand side is a different alternating sum where each term is one of the Betti numbers. Each Betti number is thus the difference between the number of cycle instigators and the non instigators one dimension higher.
Understanding the Betti Numbers
The alternating sum $\beta_{0}-\beta_{1}+\beta_{2}$ explains more about the topology of the mesh than VE+FV-E+F although the precise significance of the $\beta_{i}$ might still be a bit hazy. Below, we will try to clarify the notion of Betti numbers.
Consider a mesh with just a single vertex. That vertex is the sole entity of the mesh and a cycle instigator. It is clear that $\beta_{0}=1-0=1$ and $\beta_{1}=\beta_{2}=0$. Add another vertex, and we can create cycles consisting of either vertex or both combined: the vertices are basis vectors in a space of vertex cycles and $\beta_{0}=2-0=2$ means that the dimension of the space of vertex cycles is two. Add a non-instigating edge that connects the two vertices, and we are back to the original situation, $\beta_{0}=2-1=1$ since the now connected vertices no longer represent distinct vertex cycles. In other words, adding the (non-instigating) edge collapses the dimension of the space back down to one.
Likewise, when we add an instigating edge, it is not a single cycle but a subspace in a space of edge cycles that is created. This is because the instigator defines a unique cycle that can be construed as a basis vector. If we later add a non-instigating face that fills in the cycle, the boundary of that face also represents a basis vector which we use to project out a dimension of the space, decreasing $\beta_{1}$ by 1. A geometric way to think about it is that adding the face allows the curve to “slip across it” making edge cycles indistinguishable regardless of which way they go around the face. The upshot is that $\beta_{1}$ represents the dimension of the space of edge cycles that is not matched by faces.
The analysis of $\beta_{2}$ is simple. There is only one cycle-instigating face since a polygonal mesh is a face cycle only when it is watertight. When this face is added, $\beta_{2}=1$. If polygonal meshes were considered solids when watertight, the next step would be to add a non-instigating solid that fills in the face cycle, reducing $\beta_{2}$ to zero. Thus, the fact that we consider polygonal meshes to be hollow explains why the Euler characteristic, $\chi$, of a topological sphere is 2: a topological sphere is a single connected component with a single shell and we add one for each of these two features.
Incremental Construction
In the more general setting of simplicial complexes, Delfinado and Edelsbrunner [delfinado1995incremental] describe an incremental algorithm for computing Betti numbers. Instead of splitting mesh elements into instigators and non-instigators up front, they rely on a so-called filtration. A filtration is an arbitrary ordering of all elements of the mesh with the condition imposed that a tetrahedron can only appear in the ordering after all its triangular faces, a face can only appear after all of its edges and an edge only after its vertices.
For each element of a mesh, $\mathcal{M}$, we will use $k$ to represent its dimension: $k=0$ for vertices, $k=1$ for edges, $k=2$ for faces, and $k=3$ for tetrahedra. We will let $\mathcal{M}_{i}$ denote the mesh where all elements up to and including the $i$-th element have been added.
$\mathcal{M}$ is now built by adding the elements of the filtration one at a time. For each element, the $k$-th Betti number where $k$ is the dimension of the element, is increased by one if it is an instigator (in our terminology) with respect to $\mathcal{M}_{i}$. If it is a non-instigator, the (k1)(k-1)-th Betti number is decreased by one. Although the elements are added and subtracted one at a time, this leads to the same result as Equations (18), (19), and (20). However, it is necessary to check after each insertion whether a cycle has been created.
6 Examples
With the formulas that we have derived, it is straightforward to compute the quantitative topological properties of polygonal meshes. Figure 9 shows four shapes which are topologically quite different.
Figure 9: From left to right this image shows a hexagon (topological disc), annulus, tetrahedron (topological sphere), and torus. While all of these objects are simple they have different topology.
To find $s$, $g$, and $b$ we follow the procedure outlined in Section 4.0.1 and the Betti numbers are simply computed using Equations (18), (19), and (20). The quantitative properties of the four example shapes are summarized in Table 2.
The two triples of numbers usually do not agree, except that $\beta_{0}=s$, but since $V-E+F=2(s-g)-b=\beta_{0}-\beta_{1}+\beta_{2}$ both of these quantitative descriptions of the topology of an object do contain the same information; indeed it seems as if the information is already contained in the numbers $V$, $E$, and $F$.
Of course, we must be a little careful. Imagine an object which is a single edge shared by three trianglesmargin: [Uncaptioned image] . This object consists of 5 vertices, 7 edges, and 3 faces. We cannot compute the number of boundary curves, since they do not form closed cycles. We can compute the Betti numbers, though, and get $\beta_{0}=1$, $\beta_{1}=0$, and $\beta_{2}=0$. In other words, while this object is non-manifold it has the Betti numbers of a topological disc. Likewise, we could turn the annulus into a Möbius band and the torus into a Klein bottle – all without changing their respective Betti numbers since these objects have the same numbers of vertices, edges, and faces.
[…]
Table 2: Example meshes for which we have computed the numbers of non-instigating and instigating elements, $s$, $g$, $b$, and $\beta_{i}$.
Mesh $V_{N}$ $V_{C}$ $E_{N}$ $E_{C}$ $F_{N}$ $F_{C}$ $s$ $g$ $b$ $\beta_{0}$ $\beta_{1}$ $\beta_{2}$
Hexagon 0 6 5 1 1 0 1 0 1 1 0 0
Annulus 0 12 11 7 6 0 1 0 2 1 1 0
Tetrahedron 0 4 3 3 3 1 1 0 0 1 0 1
Torus 0 9 8 10 8 1 1 1 0 1 2 1
7 Turning Polygonal Meshes into Discs
For some applications it is necessary to transform a mesh into a topological disc. If the mesh happens to be a topological sphere a single open cut suffices since we only need to “puncture” the sphere in order to turn it into a topological disc.
For shapes of higher genus, i.e. $g>0$, it turns out that we need $2g$ closed cuts. Recall that for an object that consists of a single component and which has no boundary curves $\beta_{1}=2g$ and, for simplicity, we will assume that the object has no boundary curves. We will also assume there is just a single connected component. This means that the number of cuts we need to make is exactly the number of unique edge cycles which do not bound sets of faces. It turns out that we also need to make the required cuts exactly along these edge cycles.
In the following, we will present an algorithm originally due to Eppstein [eppstein2003dynamic], improved by Erickson and Whittlesey [erickson2005greedy] and discussed by Keenan Crane [crane_loops].
Figure 10: Left: A torus which has been cut and flattened to a disc. Note that the edges on the top are identical to those on the bottom and likewise, the left and right edges are identical. The edges of the shortest-path tree are shown pale blue, and the co-tree edges are indicated by bold arrows. In both cases, edges point from parent to child. The root of the shortest-path tree is the top left corner, and the root of the co-tree is the face indicated by a yellow star. The edges highlighted by red and purple are the cycle-instigating edges. Right: a rendering of the torus as a 3D object with the image on the left shown as a texture.
Shortest-Path Tree Construction
[…]
The first step is to create a spanning tree of the vertices. We will do this using Dijkstra’s algorithm [cormen2022introduction] which produces a tree of shortest paths from an initial vertex, $v_{0}$: the path from any vertex $v$ to $v_{0}$ along the tree is a shortest path between $v$ and $v_{0}$. Just like any other spanning tree, the shortest-path tree contains $E_{N}=V-1$ edges. A shortest-path tree computed for a torus is shown in Figure 10.
Co-Tree Construction
The next step is to construct a so-called co-tree. The co-tree is built from the dual of the mesh: a graph whose vertices are the faces of the mesh and two dual vertices are connected if the faces share an edge. We can build a co-tree using an algorithm that is similar to Dijkstra’s. The key component is a queue that keeps track of faces which have not yet been visited but are neighbours of visited faces.
An arbitrary starting face, $f_{0}$, is added to a queue. In each iteration, we pick a face $f$ from the queue. If $f$ is not in the set of visited faces, we add it to this set and then we consider each of its neighbors, i.e. each face, $f_{n}$, which shares an edge with $f$. If the edge shared by $f$ and $f_{n}$ does not belong to the shortest-path tree, we add $f_{n}$ to the queue and mark the shared edge as visited. This process is iterated until the queue is empty. A co-tree for a torus is shown in Figure 10.
Observe that the set of traversed edges must contain F1F-1 edges since the co-tree is a spanning tree over the faces. If we subtract the edges of the shortest-path tree from the total set of edges, we are down to $E_{C}=E-E_{N}$ edges. Now, we subtract $F_{N}=F-1$ edges, we have precisely $E_{C}-F_{N}=\beta_{1}$ edges left.
Simplistic Cutting Procedure
Assume now that we were to cut every edge in the shortest-path tree. The result would not have disc topology yet, but if we were to also cut the remaining $\beta_{1}$ edges, we would be left with a mesh where the only uncut edges were those that belong to the co-tree. Since a tree is connected and has no loops, this mesh would have disc topology.
Actual Cutting Procedure
However, we do not have to cut every edge of the shortest-path tree. What we do instead is to trace back from the end points of each of the $\beta_{1}$ cycle-instigating edges along the edges of the shortest-path tree. That produces $\beta_{1}$ cycles. Cutting along these edge cycles cannot disconnect the faces of the mesh since we are cutting along a subset of the edges of the simplistic procedure discussed above. However, there are no more cycles left which do not bound regions of the mesh. Consequently, these cuts must produce a mesh of disc topology. For a torus, $\beta_{1}=2$ and the two instigator edges are shown in purple and red in Figure 10. Note that in case of the red edge, the right end point is already connected to the root of the shortest path tree; in order to close the loop, we just have to trace back along the horizontal edges to the root vertex (in the top left corner). Likewise, to close the loop defined by the purple instigator, we simply trace vertically from the top vertex of the instigator to the root. It is clear that cutting along these two curves will turn the torus into a rectangle. Note that the top and bottom edges are really the same edges, and the same is true of the right side and left side edges.
Borrowing a term from homology, we will refer to the edge-cycles as generator curves. This is mostly because we need an ambiguous term to describe the curves that must be cut in order to turn the mesh into one that has disc topology. They are generators, though, in the sense that variations of these curves can be used to generate the entire group $H_{1}$.
7.1 Example
With the algorithm just described, we can compute generator curves and cut the mesh along these curves, turning it into a topological disc.
[…]
Figure 11 shows the Fertility model together with a starting vertex (orange) and eight pictures that show the individual generator curves. In Figure 12 we see the Fertility model being flattened into a disc after it has been cut along a set of generator curves
Figure 11: The fertility model and a selected vertex (top left). The remaining eight images show the generator curves (on top of a transparent rendering of the mesh) produced with this vertex as the starting point for the shortest-path tree.
Figure 12: The fertility model is here cut along each of the eight generator curves and subsequently flattened.
Note that this example was generated using the GEL library (https://github.com/janba/GEL). It is available in GEL as a Python example that can be run using PyGEL (https://pypi.org/project/PyGEL3D/).